Please help me find the least common multiple of these:
1) a^2-2ab+b^2
2) a^2-b^2
3) 3a+3b



Answer by stanbon(75887)

Please help me find the least common multiple of these:
1) a^2-2ab+b^2 = (a+b)^2
2) a^2-b^2 = (a+b)(a-b)
3) 3a+3b = 3(a+b)
-----------------
The lcm must contain each of these factors to their highest power, as follows:
lcm = 3(a-b)(a+b)^2
============Cheers,
Stan H.